### v1

@@ 3.5.1 Overview
Pre Sampling [230] [231] [232] Implements statistical parity through balanced data sampling and removes disparate impact by modifying the distribution of protected attributes in training data. Key methods include stratified sampling and balanced mini-batch creation. Safety(+), transparency(+)
Relabeling [233] [234] [235] Corrects biased labels in training data through KNN-based label propagation and gradient descent optimization. Focuses on identifying and rectifying mislabeled instances that contribute to discriminatory outcomes. Transparency(+), environmental- sustainability(+)
Represen-tation [236] [237] [238] [239] [240] Transforms raw data into fair representations through dimensionality reduction and feature engineering. Employs convex optimization to learn representations that maximize task performance while minimizing correlation with protected attributes. Transparency(+), traceability(+), safety(+)
In Treatment-driven [241] [242] [243] [244] [245] [246] [247] [248] [249] [250] Enforces equal treatment by modifying model architecture and loss functions. Incorporates fairness constraints directly into optimization objectives through adversarial training and regularization terms that penalize discriminatory decisions. Safety(+), environment- sustainability(-), transparency(+), traceability(+)
Impact-driven [251] [252] [253] [254] [255] Focuses on equalizing model outcomes across protected groups through adversarial debiasing and demographic parity constraints. Explicitly optimizes for balanced prediction distributions regardless of sensitive attributes. Safety(+), environment- sustainability(-)
Hybrid [256] [257] [258] Combines treatment and impact-driven approaches through multi-objective optimization. Uses disentangled representations and counterfactual fairness to simultaneously address disparate treatment and impact. Transparency(+), safety(+)
Post Input Correction [259] [260] Adjusts test inputs through targeted perturbation and feature transformation to ensure fair predictions. Implements systematic modifications to input data while preserving task-relevant information. Environmental- sustainability(-), transparency(+)
Classifier Correction [231] [261] [262] Modifies trained model’s decision boundaries through post-hoc calibration and threshold adjustment. Implements fairness constraints while maintaining model performance through minimal architectural changes. Environmental- sustainability(-), safety(+)
Output Correction [263] [264] Applies post-processing to model predictions using group-specific thresholds and rejection sampling. Ensures statistical parity in final outputs while maintaining prediction quality. Environmental- sustainability(-), transparency(+)
@@
-Non-discrimination in the AI Act: In the EU AI Act, the concept of non-discrimination, also refer as fairness, is not explicitly defined in a single section, but rather, it permeates various aspects of the regulation through provisions aimed at ensuring that AI systems do not create or perpetuate discrimination or bias. Firstly, the Act emphasizes the prevention of discrimination by requiring that high-risk AI systems undergo rigorous assessment processes to ensure they do not produce biased outcomes [21]. This includes testing, validation, and documentation to demonstrate that these systems can handle data fairly without leading to discriminatory results. Secondly, non-discrimination is also promoted through transparency and explainability requirements, particularly for high-risk AI systems. The Act mandates that operators provide clear information on AI systems’ functioning, capabilities, and decision-making processes [265]. This transparency helps stakeholders understand how decisions are made, thereby promoting non-discrimination in the operation of AI systems. In addition, proper management of data used by AI systems is critical to ensuring non-discrimination. The Act requires high-quality data governance practices to prevent biases arising from data misuse or poor quality. This includes measures for the accuracy, reliability, and representativeness of the data sets used, which directly influence the non-discrimination of the AI system’s outcomes. Lastly, the regulation explicitly prohibits AI practices that could manipulate or exploit vulnerable groups or otherwise lead to unfair outcomes. For example, AI systems designed to exploit vulnerabilities of individuals based on age, economic situation, or disabilities are forbidden. By integrating these principles, the AI Act seeks to ensure that AI systems used within the EU contribute to equitable outcomes, do not reinforce unfair biases, and respect the principle of non-discrimination as laid out in the broader framework of EU law.
@@
-Non-discrimination-driven approaches to public safety: Safety emphasizes the prevention of harm to individuals or data breaches due to technical failures or design flaws in AI systems, including avoiding unfair risks to specific groups due to bias. The interaction between non-discrimination and safety ensures that AI systems are safe and fair for all users. Below are particular algorithms and research examples illustrating this tension. For instance, striving for demographic parity through equal male and female parole rates may inadvertently disadvantage lower-risk female inmates to fulfill this ratio, thereby breaking the non-discrimination principle of equalized odds [266]. The application of machine learning in critical sectors such as the judiciary, welfare systems, and autonomous driving underscores the myriad ways in which AI systems, imbued with inherent biases, can impact daily life and the subtleties of biases in AI and robotics infiltrating real-world scenarios [267], emphasizing the non-discrimination imperative for researchers and engineers to anticipate downstream effects in AI system development. Acknowledging the predictive value of gender in such contexts could lead to a decision-making paradigm that is unfair to minority groups, without secured public safety outcomes. By implementing these approaches, AI systems can contribute to public safety through equitable law enforcement, unbiased emergency response, inclusive public health measures, and fair access to public services. Ultimately, non-discriminatory AI helps reduce social tensions, improve trust in institutions, and ensure that technological advancements benefit all members of society equally [268].
@@ 3.5.2 Non-discrimination versus transparency, traceability, safety, and environmental sustainability
Transparency and explanation [269][270] are fundamental in fostering trust and understanding in machine learning non-discrimination. By identifying and addressing the underlying causes of bias, causal methods can help reveal underlying biases and improve decision-making transparency. While Inverse Propensity Scoring (IPS) techniques [271] offer a useful approach for enhancing fairness in AI systems by tackling selection bias and promoting more equitable datasets, they are not without drawbacks. These methods, though practical, struggle to adapt to changes in observational patterns and are frequently designed for particular use cases rather than being universally applicable.By introducing small changes to input data, perturbation methods can reveal hidden biases in AI models, helping identify areas where discrimination may occur. Studies using perturbation techniques have demonstrated that implementing measures to ensure non-discrimination does not substantially reduce accuracy in AI systems [272][273]. This finding supports the use of such interventions to enhance privacy protection, transparency and fairness in AI applications. Interpretable models [274] further non-discrimination by making decision processes transparent, advocating for a cohesive approach to creating fair and transparent algorithmic systems [275].
@@
Studies [237] introduce a pre-processing strategy to mitigate bias by altering or removing bias-originating documents during training, addressing word embedding bias. Concurrently, research [276] develops an encryption method for sensitive user data, enhancing security and privacy while allowing non-discrimination verification without exposing data to unauthorized use or access.
The impact of non-discrimination AI approaches to sustainable AI: Repairing or rebuilding a fair model is typically time-consuming, particularly when post-processing is required [264, 231, 259]. Adversarial learning techniques, such as adversarial debiasing, eliminate discrimination using a discriminator, but introducing additional model training and inference processes is not environmentally friendly [251, 244, 255]. However, some repair methods are relatively more sustainable than rebuilding. For example, using a suitable algorithm to learn counterfactual distributions can repair a black-box classifier without the need for retraining, offering an environmentally sustainable way to restore the model’s fairness [235].
@@ 4.1 AI Act’s technical challenges for human-centric design
Strengthening AI systems against supply-chain vulnerabilities: Addressing supply-chain vulnerabilities in AI systems extends beyond protecting against direct attacks to include securing the entire ecosystem surrounding AI development, from the data sources and software libraries to the deployment environments. These vulnerabilities may arise from compromised data leading to poisoned models, tampered software libraries inducing backdoor threats, or adversarial manipulations undermining model integrity, availability, and privacy. Effective mitigation requires a comprehensive approach that combines advanced technical strategies, like adversarial training for resilience against attacks [128], [222], and rigorous data sanitization [77] to prevent malicious data injection, with robust procedural and policy frameworks. These measures must account for the inherent trade-offs between model accuracy and robustness, and the dynamic nature of AI threats necessitates continuous vigilance and adaptation. Enhancing the security of AI systems against supply-chain vulnerabilities thus involves a concerted effort across the AI community to foster secure, reliable, and trustworthy AI through ongoing collaboration, innovation, and a strong commitment to security best practices.
@@
The dual-edged nature of technologies in ethical AI promotion: Promoting ethical AI underscores the nuanced balance between harnessing innovative technological advances to foster ethical AI practices and navigating the inherent challenges these technologies may pose. On one side, technologies like machine learning algorithms, blockchain, and data analytics tools can significantly enhance transparency, accountability, and fairness in AI systems, laying the foundation for more ethical AI development. They provide mechanisms for unbiased decision-making, environmental friendliness, secure data sharing, and enhanced privacy, which are crucial for ethical standards. For example, transfer learning, as discussed in Section 3.3.2, typically has a positive influence on the development of AI models, particularly in terms of environmental sustainability, non-discrimination, and transparency. Nonetheless, not all transfer learning approaches yield beneficial outcomes for environmentally sustainable development, and some are necessary only in specific scenarios. For example, instance-based transfer learning often has a detrimental environmental impact, while feature-based transfer learning might have a slightly negative effect. In situations where the training set is small, and label information originates solely from the source domain, users might be constrained to opt for either instance-based or feature-based transfer learning over more eco-friendly alternatives. According to [191], other environmental sustainability transfer learning methods can be chosen when the label information of the target-domain instances is available. Otherwise, a trade-off between environmental sustainability and performance shall be considered [175]. Additionally, the efficiency of instance-based or feature-based transfer learning in enhancing non-discrimination and transparency while remaining cost-effective is yet to be determined.
@@
Another example is adversarial training, as discussed in Section 3.1.2 and Section 3.5.2. It significantly enhances the safety and non-discrimination aspects of AI model development, respectively [66, 67, 241, 244]. This method strengthens the model’s resilience or reduces its dependency on certain specific features by introducing adversarial examples during the training phase. However, while aimed at promoting model fairness, adversarial training might inadvertently compromise the model’s performance in critical scenarios, which leads to potential threats in safety [244], such as identifying illegal weapon possession [241]. Adversarial training with crafted adversarial examples introduces the features that the model learns to treat independently, enhancing the system’s transparency and traceability by offering clear insights into the model’s robustness/defenses against particular adversarial inputs [66, 67]. Further, adversarial training may introduce a more complex training process, resulting in more computational cost and potentially conflicting with environmental sustainability objectives [241].
@@
Differential privacy (DP) is the third technique discussed due to its dual-edged nature. Generally, DP enhances AI transparency and safety but poses challenges for developing environmentally friendly AI and complicates traceability, with a complicated effect on fairness. This technique is adaptable and applicable across various stages of AI development, including training data preprocessing, feature engineering, model training, and prediction, primarily to ensure privacy-preserving data analysis [84, 277]. Such applications help AI providers establish suitable levels of transparency tailored for different users or customers. In addition, with DP techniques, the AI model can be more robust and safe by mitigating various attacks like membership inference attacks and model inversion attacks [278, 279]. However, incorporating DP increases the complexity of models, learning tasks, and datasets, particularly during the model training phase, which can contradict the principles of environmental sustainability [278]. As for traceability, although the DP-based model is more complex to understand than the typical model, the DP-based explainable model shows more private interpretability. Patel et al. [280] proposed an optimization method to minimize the total privacy loss while maintaining a high explanation quality. The impact of DP on the fairness of AI systems is complex. On the one hand, it supports fairness by enabling re-sampling of training data [277], which helps maintain balance across different groups. On the other hand, using differentially private training data can inadvertently introduce biases, potentially affecting fairness negatively [281], though these effects might be mitigated by perturbing the model’s outputs [282]. Thus, while DP contributes positively to transparency and safety, its influence on environmental sustainability, traceability, and fairness in AI necessitates careful consideration and balanced application.
@@
Ensuring Ethical Development in Large Language Models: The importance of data in the development of Large Language Models (LLMs) cannot be overstated. The correlation between the growth of these models and the exponential increase in required training data is evident in their evolution. Although detailed disclosure of data sources by LLMs remains rare, the instances where such transparency exists reveal the vast data consumption involved in training [283]. The emergence of chatbots [284] marks a significant technological leap, offering potential across a wide range of business applications, from entertainment to critical sectors. However, this advancement comes with challenges, notably the vulnerability of LLMs to prompt injections [107], a technique that can elicit undesired responses. Recent study has identified inherent limitations in strictly censoring LLMs, highlighting the need for alternative risk management strategies such as implementing controlled access points for models and enhancing cybersecurity measures[285]. Despite the impressive capabilities demonstrated by chatbots in specific tasks, the technology is still in its infancy and requires cautious deployment, especially in scenarios where trustworthiness is paramount. The need for continuous monitoring underscores the technology’s emerging status and the inherent challenges in ensuring ethical AI development. Future directions should focus on increasing transparency around data sources, enhancing security measures to counter vulnerabilities, and adhering to rigorous ethical standards to foster trust and reliability in LLM applications.
@@
Trade-offs between the attributes of ethical and trustworthy AI: To navigate towards the goal of ethical AI, it is paramount to recognize and address the intricate balance among the diverse attributes that underpin the trustworthiness of AI systems. These attributes include human-centric principles that are defined in regulation as well as some that extend beyond, such as accuracy, susceptibility to adversarial attacks, explainability, privacy, and robustness to adversarial attacks. A singular focus on optimizing one attribute, such as accuracy, often results in compromises in other critical areas like fairness and vulnerability to attacks [286]. This delicate balancing act between various desirable qualities highlights a key challenge in AI development: the inherent trade-offs that must be managed. For instance, enhancing an AI system’s adversarial robustness might inadvertently lower its accuracy and fairness [287].
@@
The path forward involves a multifaceted approach that does not seek to maximize performance in one attribute at the expense of others but rather aims for a harmonious balance that upholds the principles of AI regulation. This requires a comprehensive understanding of the interplay between different AI system characteristics, necessitating ongoing research into characterizing and navigating the trade-offs between them. As AI technology becomes increasingly integrated into various facets of modern life, the importance of such research grows, underscoring the need for innovative solutions that ensure AI systems are trustworthy by being fair, secure, transparent, traceable, and environmentally friendly. Future directions in achieving ethical AI must focus on developing methodologies and technologies that enable this balance, ensuring AI systems are designed and deployed in a manner that earns and maintains public trust [27].

### v2

@@ 3.5.1 Overview
Pre Sampling (?, ?, ?) Implements statistical parity through balanced data sampling and removes disparate impact by modifying the distribution of protected attributes in training data. Key methods include stratified sampling and balanced mini-batch creation. Safety (+), transparency (+)
Relabeling (?, ?, ?) Corrects biased labels in training data through KNN-based label propagation and gradient descent optimization. Focuses on identifying and rectifying mislabeled instances that contribute to discriminatory outcomes. Transparency (+), environmental-sustainability (+)
Represen-tation (?, ?, ?, ?) Transforms raw data into fair representations through dimensionality reduction and feature engineering. Employs convex optimization to learn representations that maximize task performance while minimizing correlation with protected attributes. Transparency (+), traceability (+), safety (+)
In Treatment-driven (?, ?, ?, ?, ?) Enforces equal treatment by modifying model architecture and loss functions. Incorporates fairness constraints directly into optimization objectives through adversarial training and regularization terms that penalize discriminatory decisions. Safety (+), environment-sustainability (-), transparency(+), traceability (+)
Impact-driven (?, ?, ?, ?, ?) Focuses on equalizing model outcomes across protected groups through adversarial debiasing and demographic parity constraints. Explicitly optimizes for balanced prediction distributions regardless of sensitive attributes. Safety (+), environment-sustainability (-)
Hybrid (?, ?, ?) Combines treatment and impact-driven approaches through multi-objective optimization. Uses disentangled representations and counterfactual fairness to simultaneously address disparate treatment and impact. Transparency (+), safety (+)
Post Input correction (?, ?) Adjusts test inputs through targeted perturbation and feature transformation to ensure fair predictions. Implements systematic modifications to input data while preserving task-relevant information. Environmental-sustainability (-), transparency (+)
Classifier correction (?, ?, ?) Modifies trained model’s decision boundaries through post-hoc calibration and threshold adjustment. Implements fairness constraints while maintaining model performance through minimal architectural changes. Environmental-sustainability (-), safety (+)
Output correction (?, ?) Applies post-processing to model predictions using group-specific thresholds and rejection sampling. Ensures statistical parity in final outputs while maintaining prediction quality. Environmental-sustainability (-), transparency (+)
@@
-Non-discrimination in the AI Act: In the EU AI Act, the concept of non-discrimination, also refer as fairness, is not explicitly defined in a single section, but rather, it permeates various aspects of the regulation through provisions aimed at ensuring that AI systems do not create or perpetuate discrimination or bias. Firstly, the Act emphasizes the prevention of discrimination by requiring that high-risk AI systems undergo rigorous assessment processes to ensure they do not produce biased outcomes (?). This includes testing, validation, and documentation to demonstrate that these systems can handle data fairly without leading to discriminatory results. Secondly, non-discrimination is also promoted through transparency and explainability requirements, particularly for high-risk AI systems. The Act mandates that operators provide clear information on AI systems’ functioning, capabilities, and decision-making processes (?). This transparency helps stakeholders understand how decisions are made, thereby promoting non-discrimination in the operation of AI systems. In addition, proper management of data used by AI systems is critical to ensuring non-discrimination. The Act requires high-quality data governance practices to prevent biases arising from data misuse or poor quality. This includes measures for the accuracy, reliability, and representativeness of the data sets used, which directly influence the non-discrimination of the AI system’s outcomes. Lastly, the regulation explicitly prohibits AI practices that could manipulate or exploit vulnerable groups or otherwise lead to unfair outcomes. For example, AI systems designed to exploit vulnerabilities of individuals based on age, economic situation, or disabilities are forbidden. By integrating these principles, the AI Act seeks to ensure that AI systems used within the EU contribute to equitable outcomes, do not reinforce unfair biases, and respect the principle of non-discrimination as laid out in the broader framework of EU law.
@@
-Non-discrimination-driven approaches to public safety: Safety emphasizes the prevention of harm to individuals or data breaches due to technical failures or design flaws in AI systems, including avoiding unfair risks to specific groups due to bias. The interaction between non-discrimination and safety ensures that AI systems are safe and fair for all users. Below are particular algorithms and research examples illustrating this tension. For instance, striving for demographic parity through equal male and female parole rates may inadvertently disadvantage lower-risk female inmates to fulfill this ratio, thereby breaking the non-discrimination principle of equalized odds (?). The application of machine learning in critical sectors such as the judiciary, welfare systems, and autonomous driving underscores the myriad ways in which AI systems, imbued with inherent biases, can impact daily life and the subtleties of biases in AI and robotics infiltrating real-world scenarios (?), emphasizing the non-discrimination imperative for researchers and engineers to anticipate downstream effects in AI system development. Acknowledging the predictive value of gender in such contexts could lead to a decision-making paradigm that is unfair to minority groups, without secured public safety outcomes. By implementing these approaches, AI systems can contribute to public safety through equitable law enforcement, unbiased emergency response, inclusive public health measures, and fair access to public services. Ultimately, non-discriminatory AI helps reduce social tensions, improve trust in institutions, and ensure that technological advancements benefit all members of society equally (?).
@@ 3.5.2 Non-discrimination versus transparency, traceability, safety, and environmental sustainability
Transparency and explanation (?)(?) are fundamental in fostering trust and understanding in machine learning non-discrimination. By identifying and addressing the underlying causes of bias, causal methods can help reveal underlying biases and improve decision-making transparency. While Inverse Propensity Scoring (IPS) techniques (?) offer a useful approach for enhancing fairness in AI systems by tackling selection bias and promoting more equitable datasets, they are not without drawbacks. These methods, though practical, struggle to adapt to changes in observational patterns and are frequently designed for particular use cases rather than being universally applicable.By introducing small changes to input data, perturbation methods can reveal hidden biases in AI models, helping identify areas where discrimination may occur. Studies using perturbation techniques have demonstrated that implementing measures to ensure non-discrimination does not substantially reduce accuracy in AI systems (?)(?). This finding supports the use of such interventions to enhance privacy protection, transparency and fairness in AI applications. Interpretable models (?) further non-discrimination by making decision processes transparent, advocating for a cohesive approach to creating fair and transparent algorithmic systems (?).
@@
Studies (?) introduce a pre-processing strategy to mitigate bias by altering or removing bias-originating documents during training, addressing word embedding bias. Concurrently, research (?) develops an encryption method for sensitive user data, enhancing security and privacy while allowing non-discrimination verification without exposing data to unauthorized use or access.
The impact of non-discrimination AI approaches to sustainable AI: Repairing or rebuilding a fair model is typically time-consuming, particularly when post-processing is required (?, ?, ?). Adversarial learning techniques, such as adversarial debiasing, eliminate discrimination using a discriminator, but introducing additional model training and inference processes is not environmentally friendly (?, ?, ?). However, some repair methods are relatively more sustainable than rebuilding. For example, using a suitable algorithm to learn counterfactual distributions can repair a black-box classifier without the need for retraining, offering an environmentally sustainable way to restore the model’s fairness (?).
@@ 4.1 AI Act’s technical challenges for human-centric design
Strengthening AI systems against supply-chain vulnerabilities: Addressing supply-chain vulnerabilities in AI systems extends beyond protecting against direct attacks to include securing the entire ecosystem surrounding AI development, from the data sources and software libraries to the deployment environments. These vulnerabilities may arise from compromised data leading to poisoned models, tampered software libraries inducing backdoor threats, or adversarial manipulations undermining model integrity, availability, and privacy. Effective mitigation requires a comprehensive approach that combines advanced technical strategies, like adversarial training for resilience against attacks (?), (?), and rigorous data sanitization (?) to prevent malicious data injection, with robust procedural and policy frameworks. These measures must account for the inherent trade-offs between model accuracy and robustness, and the dynamic nature of AI threats necessitates continuous vigilance and adaptation. Enhancing the security of AI systems against supply-chain vulnerabilities thus involves a concerted effort across the AI community to foster secure, reliable, and trustworthy AI through ongoing collaboration, innovation, and a strong commitment to security best practices.
@@
The dual-edged nature of technologies in ethical AI promotion: Promoting ethical AI underscores the nuanced balance between harnessing innovative technological advances to foster ethical AI practices and navigating the inherent challenges these technologies may pose. On one side, technologies like machine learning algorithms, blockchain, and data analytics tools can significantly enhance transparency, accountability, and fairness in AI systems, laying the foundation for more ethical AI development. They provide mechanisms for unbiased decision-making, environmental friendliness, secure data sharing, and enhanced privacy, which are crucial for ethical standards. For example, transfer learning, as discussed in Section 3.3.2, typically has a positive influence on the development of AI models, particularly in terms of environmental sustainability, non-discrimination, and transparency. Nonetheless, not all transfer learning approaches yield beneficial outcomes for environmentally sustainable development, and some are necessary only in specific scenarios. For example, instance-based transfer learning often has a detrimental environmental impact, while feature-based transfer learning might have a slightly negative effect. In situations where the training set is small, and label information originates solely from the source domain, users might be constrained to opt for either instance-based or feature-based transfer learning over more eco-friendly alternatives. According to (?), other environmental sustainability transfer learning methods can be chosen when the label information of the target-domain instances is available. Otherwise, a trade-off between environmental sustainability and performance shall be considered (?). Additionally, the efficiency of instance-based or feature-based transfer learning in enhancing non-discrimination and transparency while remaining cost-effective is yet to be determined.
@@
Another example is adversarial training, as discussed in Section 3.1.2 and Section 3.5.2. It significantly enhances the safety and non-discrimination aspects of AI model development, respectively (?, ?, ?, ?). This method strengthens the model’s resilience or reduces its dependency on certain specific features by introducing adversarial examples during the training phase. However, while aimed at promoting model fairness, adversarial training might inadvertently compromise the model’s performance in critical scenarios, which leads to potential threats in safety (?), such as identifying illegal weapon possession (?). Adversarial training with crafted adversarial examples introduces the features that the model learns to treat independently, enhancing the system’s transparency and traceability by offering clear insights into the model’s robustness/defenses against particular adversarial inputs (?, ?). Further, adversarial training may introduce a more complex training process, resulting in more computational cost and potentially conflicting with environmental sustainability objectives (?).
@@
Differential privacy (DP) is the third technique discussed due to its dual-edged nature. Generally, DP enhances AI transparency and safety but poses challenges for developing environmentally friendly AI and complicates traceability, with a complicated effect on fairness. This technique is adaptable and applicable across various stages of AI development, including training data preprocessing, feature engineering, model training, and prediction, primarily to ensure privacy-preserving data analysis (?, ?). Such applications help AI providers establish suitable levels of transparency tailored for different users or customers. In addition, with DP techniques, the AI model can be more robust and safe by mitigating various attacks like membership inference attacks and model inversion attacks (?, ?). However, incorporating DP increases the complexity of models, learning tasks, and datasets, particularly during the model training phase, which can contradict the principles of environmental sustainability (?). As for traceability, although the DP-based model is more complex to understand than the typical model, the DP-based explainable model shows more private interpretability. Patel et al. (?) proposed an optimization method to minimize the total privacy loss while maintaining a high explanation quality. The impact of DP on the fairness of AI systems is complex. On the one hand, it supports fairness by enabling re-sampling of training data (?), which helps maintain balance across different groups. On the other hand, using differentially private training data can inadvertently introduce biases, potentially affecting fairness negatively (?), though these effects might be mitigated by perturbing the model’s outputs (?). Thus, while DP contributes positively to transparency and safety, its influence on environmental sustainability, traceability, and fairness in AI necessitates careful consideration and balanced application.
@@
Ensuring Ethical Development in Large Language Models: The importance of data in the development of Large Language Models (LLMs) cannot be overstated. The correlation between the growth of these models and the exponential increase in required training data is evident in their evolution. Although detailed disclosure of data sources by LLMs remains rare, the instances where such transparency exists reveal the vast data consumption involved in training (?). The emergence of chatbots (?) marks a significant technological leap, offering potential across a wide range of business applications, from entertainment to critical sectors. However, this advancement comes with challenges, notably the vulnerability of LLMs to prompt injections (?), a technique that can elicit undesired responses. Recent study has identified inherent limitations in strictly censoring LLMs, highlighting the need for alternative risk management strategies such as implementing controlled access points for models and enhancing cybersecurity measures(?). Despite the impressive capabilities demonstrated by chatbots in specific tasks, the technology is still in its infancy and requires cautious deployment, especially in scenarios where trustworthiness is paramount. The need for continuous monitoring underscores the technology’s emerging status and the inherent challenges in ensuring ethical AI development. Future directions should focus on increasing transparency around data sources, enhancing security measures to counter vulnerabilities, and adhering to rigorous ethical standards to foster trust and reliability in LLM applications.
@@
Trade-offs between the attributes of ethical and trustworthy AI: To navigate towards the goal of ethical AI, it is paramount to recognize and address the intricate balance among the diverse attributes that underpin the trustworthiness of AI systems. These attributes include human-centric principles that are defined in regulation as well as some that extend beyond, such as accuracy, susceptibility to adversarial attacks, explainability, privacy, and robustness to adversarial attacks. A singular focus on optimizing one attribute, such as accuracy, often results in compromises in other critical areas like fairness and vulnerability to attacks (?). This delicate balancing act between various desirable qualities highlights a key challenge in AI development: the inherent trade-offs that must be managed. For instance, enhancing an AI system’s adversarial robustness might inadvertently lower its accuracy and fairness (?).
@@
The path forward involves a multifaceted approach that does not seek to maximize performance in one attribute at the expense of others but rather aims for a harmonious balance that upholds the principles of AI regulation. This requires a comprehensive understanding of the interplay between different AI system characteristics, necessitating ongoing research into characterizing and navigating the trade-offs between them. As AI technology becomes increasingly integrated into various facets of modern life, the importance of such research grows, underscoring the need for innovative solutions that ensure AI systems are trustworthy by being fair, secure, transparent, traceable, and environmentally friendly. Future directions in achieving ethical AI must focus on developing methodologies and technologies that enable this balance, ensuring AI systems are designed and deployed in a manner that earns and maintains public trust (?).